a triangle has an area of 90 square centimeters and a height of 12 centimeters.wath is the length of the base
Answer by Boreal(15235)

Area = (1/2) bh; Area=90 cm^2, height=12 cm base is x cm.
90 cm^2= (1/2) *12 cm *x cm Units are fine. Multiply both sides by 2 to remove the fraction (1/2)
180=12 * base
180/12 = base=15 cm.